OpenStudy (anonymous):

Andrew claims that a wooden bookend in the shape of a 45°- 45°- 90° right triangle has side lengths of 9 in., 9 in., and 11 in. Is he correct? If so, show the work and if not, show why not.

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

He is not correct. 9^2 + 9^2 not = 11^2 because 9^2 + 9^2 = 81 + 81 = 162 and 11^2 = 121 162 not = 121

OpenStudy (anonymous):

If this were a right triangle, then we would have a^2 + b^2 = c^2, but we don't have that here.